OpenStudy (anonymous):

What is the sum of an 11–term geometric series if the first term is 4 and the last term is 4,096? 11,264 8,188 45,056 9,124

OpenStudy (anonymous):

the answer is B...8188...and the common ratio is 2

OpenStudy (anonymous):

ratio is 2 (s2 ^11 = 4096) so from http://en.wikipedia.org/wiki/Geometric_series the sun = 4(1 -2 ^11)/(1-2) = 8088